OpenStudy (anonymous):

Victoria has some money from her birthday and the amount is modeled by the function h(x) = 200. She read about a bank that has savings accounts that accrue interest according to the function s(x) = (1.05)^x - 1. Victoria is thinking about putting her money into the savings account to gain interest. Using complete sentences, explain to Victoria how she can combine her functions to create a new function, and explain what this new function means.

OpenStudy (solomonzelman):

s(x) is the function for how much percent of interest she will gain?

OpenStudy (anonymous):

I just don't know how to combine the functions like do we multiply divide add subtract???

OpenStudy (solomonzelman):

wait, so the \(\large\color{black}{ s(x) }\) is basically saying that for the first year it will be \(\large\color{black}{ s(1) }\) , for the 2nd year it will be \(\large\color{black}{ s(2) }\) , and on.... like this?

OpenStudy (anonymous):

Yes I do believe that x is the number of years...

OpenStudy (solomonzelman):

but it wouldn't be that s(2) is the combined interest after the 1st 2 years (altogether) ?

OpenStudy (solomonzelman):

I would assume that it is not so. and will see how it comes out.

OpenStudy (anonymous):

I think that if we do 2 years we would make the function Sn=200+1.05^2-1

OpenStudy (solomonzelman):

it depends in whether the entire interest for 1st 2 years would be \(\large\color{black}{s(1)+s(2) }\) , or just \(\large\color{black}{s(2) }\) . I am sure about this from the wording of the question.

OpenStudy (anonymous):

That went to 210 it kinda seems right??

OpenStudy (anonymous):

I think you would add the 200 since she got that money and then we have to add the interest that she receives

OpenStudy (solomonzelman):

you multiply. lets say your increase is + 0.04 of the sum, then you would do sum * 1.04

OpenStudy (anonymous):

I still got 210 as my answer???

OpenStudy (solomonzelman):

you are asked to make one function that would be giving you the interest based on the s(x), and with the initial sum of 200 (as in h(x) ).

OpenStudy (solomonzelman):

no that was illogical just now, \(\large\color{black}{s(x \times h) =200[~(1.05)^x - 1~] }\) would be giving you the function for the just additional money.

OpenStudy (solomonzelman):

am I interpreting the s(x) correctly, maybe x-1 is the exponent?

OpenStudy (anonymous):

\[s(x)=1.05^(x-1)\]

OpenStudy (solomonzelman):

\(\large\color{black}{s(x) =(1.05)^{x-1} }\) like this?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

oh that is better

OpenStudy (solomonzelman):

\(\large\color{black}{s(x) =(1.05)^{x-1} }\) is just a geometric sequence where: ~\(\large\color{black}{r=1.05 }\) ~\(\large\color{black}{a_1=1 }\)

OpenStudy (anonymous):

Now how do we combine them that is my problem

OpenStudy (solomonzelman):

\(\large\color{black}{s(x \times h) }\)

OpenStudy (solomonzelman):

do you know how to write \(\large\color{black}{s(x \times h) }\) if: ~ \(\large\color{black}{s( h)=200 }\) ~ \(\large\color{black}{s(x) =(1.05)^{x-1} }\)

OpenStudy (anonymous):

How would you solve that?

OpenStudy (solomonzelman):

If I had: \(\large\color{black}{s(h ) =50 }\) \(\large\color{black}{s(x) =(2.07)^{x-2} }\) then \(\large\color{black}{s(x \times h) =50(2.07)^{x-2} }\)

OpenStudy (solomonzelman):

now, if I have: \(\large\color{black}{s(h ) =200 }\) \(\large\color{black}{s(x) =(1.05)^{x-1} }\) then \(\large\color{black}{s(x \times h) =? }\)

OpenStudy (anonymous):

I don't know how to solve the exponent to get it down

OpenStudy (solomonzelman):

now, if I have: \(\large\color{black}{s(h ) =200 }\) \(\large\color{black}{s(x) =(1.05)^{x-1} }\) then \(\large\color{black}{s(x \times h) =200(1.05)^{x-1} }\)

OpenStudy (solomonzelman):

see what i am doing>?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I think I got it thanks

OpenStudy (solomonzelman):

yw